OpenStudy (anonymous):

A gas occupies 1560 cm3 at 285 K. To what temp must the gas be lowered to, if it is to occupy 25.0 cm3? Assume a constant pressure.

OpenStudy (anonymous):

would it be set up like this 285 K/1560cm3 X ?/25.0cm3

OpenStudy (aaronq):

yeah, you can set up it either way \[\frac{ T _{1} }{ V _{1} }=\frac{ T _{2} }{ V _{2} } or \frac{ V _{1} }{ T _{1} }=\frac{ V _{2} }{ T _{2} } \]

OpenStudy (anonymous):

so then it would be 444600/1560=285K is that the final answer???

OpenStudy (anonymous):

I messed up somewhere

OpenStudy (aaronq):

285/1560=x/25 -> x=4.5673076923076923 K

OpenStudy (anonymous):

0.1826K??

OpenStudy (aaronq):

thats not what i got

OpenStudy (anonymous):

so you don't take 4.57307/25.0???

OpenStudy (aaronq):

no, you're solving for the T when the gas occupies 25 cm^3, which x

OpenStudy (aaronq):

which is*

OpenStudy (anonymous):

multiply??? 4.567307 by 25.0cm3

OpenStudy (anonymous):

would i take 1560/25.0=6.24

OpenStudy (aaronq):

nonono, This is the set up: 285 K/1560cm3 =T/25.0cm3 -> T=285 K*25.0cm3/1560cm3= 4.57 K

OpenStudy (anonymous):

thank you again ;D

OpenStudy (aaronq):

hah no problem !

OpenStudy (anonymous):

I wish it was that easy for me!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!! ;D

OpenStudy (anonymous):

tHANKYOU

OpenStudy (aaronq):

haha practice is all it takes :P you're on your own now :S, though, i gotta go study got a test tmorw

OpenStudy (anonymous):

Good luck not that you'll need it

OpenStudy (anonymous):

your a life saver